Tri Star, Inc., has the following mutually exclusive projects:

Year Project A Project B
0 –$ 13,700 –$ 9,100
1 8,300 3,800
2 6,900 3,300
3 2,100 5,700


Calculate the payback period for each project.
Answer #1

The answers are:

A's payback 1.78 years
B's payback 2.35 years

Calculations are shown below:

Payback period for project A:

Year Cash flow of Project A Cumulative cash flow
0 -13,700.00 -13,700.00
1 8,300.00 -5,400.00
2 6,900.00 1,500.00
3 2,100.00

Thus payback for project A = 1 year+(5400/6900) = 1.78 years

Payback for project B:

Year Cash flow of Project B Cumulative cash flow
0 -9,100.00 -9,100.00
1 3,800.00 -5,300.00
2 3,300.00 -2,000.00
3 5,700.00 3,700.00

Payback for B = 2+(2000/5700) = 2.35 years